a bird is flying towards north with velocity 5m/s and a dog is running due east . find velocity of dog seen by the bird . given velocity of dog = 5√3m/s

Grade:12th pass

1 Answers

Utsav Basu
6 years ago
Let’s consider the velocity of bird to be 5j nd dog to be (5√3)i
So, velocity of dod seen by bird (in vector form)
= vdog – vbird
= (5√3)i – 5j
So, magnitude of v=5((√3)2 + 12)1/2=10
Direction of velocity= tan-1(-5/5√3)= -60deg.